• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Kalyan

    Log entry time 16:43:34 on February 08, 2009

    Entry number 259835

    This entry is a followup to: 259812

    keyword=BB scalers gates problem fixed

    The negative helicity gate problem on the BB is fixed. The problem was that the negative helicity state signal got attenuated to below NIM standard and the module PS757 was not firing on it anymore. So the gated were always on and scalers were counting just as though it were ungated, hence they were same as ungated(see halog 259812)

    Bob looked at it and suggested to make a copy on other module. Now the gates are working.